Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

LightBulbs

Expert replies
by ani781 » Thu Sep 26, 2013 4:12 am
string of 10 lightbulbs is wired in such a way that if
any individual lightbulb fails, the entire string fails. If for
each individual lightbulb the probability of failing during
time period T id 0.06, what is the probability that the
string of lightbulbs will fail during time period T?

A.0.06
B.(0.06)^10
C.1-(0.06)^10
D.(0.94)^10
E.1-(0.94)^10


How do I approach these questions ?
Any pointer / book or example questions of similar kind would be very very helpful.

Thanks.
Join the discussion
Source: — Problem Solving |

by theCodeToGMAT » Thu Sep 26, 2013 4:56 am
Answer would be {E}

Probability of Non Failure = 1 - 0.06 = 0.94

Probability of 10 bulbs not failing = (0.94)^10

So, probability of failure = 1 - probability of non failure = 1 - (0.94)^10

what is the OA?
R A H U L
Join the discussion

by ani781 » Thu Sep 26, 2013 5:25 am
You are really nice at these. Thanks; your answer is indeed the OA.
Join the discussion

by Brent@GMATPrepNow » Thu Sep 26, 2013 5:28 am
ani781 wrote:string of 10 lightbulbs is wired in such a way that if
any individual lightbulb fails, the entire string fails. If for
each individual lightbulb the probability of failing during
time period T id 0.06, what is the probability that the
string of lightbulbs will fail during time period T?

A.0.06
B.(0.06)^10
C.1-(0.06)^10
D.(0.94)^10
E.1-(0.94)^10

Aside: If P(bulb fails) = 0.06, then P(bulb doesn't fail) = 0.94


Okay, the entire string of lightbulbs will fail if 1 or more lightbulbs fail.
So, we want P(at least 1 lightbulb fails)

When it comes to probability questions involving "at least," it's best to try using the complement.
That is, P(Event A happening) = 1 - P(Event A not happening)
P(at least 1 lightbulb fails) = 1 - P(zero lightbulbs fail)


P(zero lightbulbs fail)
P(zero lightbulbs fail) = P(1st bulb doesn't fail AND 2nd bulb doesn't fail AND 3rd bulb doesn't fail AND . . . AND 9th bulb doesn't fail AND 10th bulb doesn't fail)
= P(1st bulb doesn't fail) x P(2nd bulb doesn't fail) x P(3rd bulb doesn't fail) x . . . x P(9th bulb doesn't fail) x P(10th bulb doesn't fail)
= (0.94) x (0.94) x (0.94) x . . . x(0.94) x (0.94)
= (0.96)^10

So, P(at least 1 lightbulb fails) = 1 - P(zero lightbulbs fail)
= 1 - (0.94)^10
= E

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by ani781 » Thu Sep 26, 2013 5:37 am
Thanks a lott Brent. Now this makes it very clear to me.

Is it possible for you to point me at some similar or almost in the same line of questions ? Or perhaps you can guide me to some more of these questions to further solidify my understanding ?


Thanks in advance.[/quote]
Join the discussion

by Brent@GMATPrepNow » Thu Sep 26, 2013 6:00 am
ani781 wrote:Thanks a lott Brent. Now this makes it very clear to me.

Is it possible for you to point me at some similar or almost in the same line of questions ? Or perhaps you can guide me to some more of these questions to further solidify my understanding ?


Thanks in advance.
[/quote]

Sure thing.
Try these:
- https://www.beatthegmat.com/how-to-solve ... 16924.html
- https://www.beatthegmat.com/probability- ... 21264.html
- https://www.beatthegmat.com/probability- ... 23109.html
- https://www.beatthegmat.com/bird-t265208.html
- https://www.beatthegmat.com/probability-t264072.html

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Gurpreet singh » Tue Jun 28, 2016 5:58 am
How do we solve it without using the complement method?
If the probability of one bulb failing is .06 is the probability of 10 bulb failing not= (.06)^10?
Join the discussion

by DavidG@VeritasPrep » Tue Jun 28, 2016 7:48 am
Gurpreet singh wrote:How do we solve it without using the complement method?
If the probability of one bulb failing is .06 is the probability of 10 bulb failing not= (.06)^10?
(.06)^10 would be the probability of all 10 bulbs failing. But the entire string fails if at least one bulb fails. So there are going to be a lot of scenarios where this can happen.

If 1 fails and 9 work, the string fails
If 2 fail and 8 work, the string fails
If 3 fail and 7 work, the string fails
etc.

It would take a very long time to calculate all of these individually. So you'd want to approach this the way Brent did, and solve for 1 - P(no bulbs fail)
Veritas Prep | GMAT Instructor

Veritas Prep Reviews
Save $100 off any live Veritas Prep GMAT Course
Join the discussion

by Gurpreet singh » Tue Jun 28, 2016 4:29 pm
Thank you.
Join the discussion